Your search - a turntable is a uniform disk of mass 12kg and radius 0.13metres. the turntable is spinning at initial constant frequency of 33 cycles per minute. the motor is turned off and the turntable slows down to a stop in 8 seconds due to frictional torque. what is the magnitude of the frictional torque acting on the table -


1
Expert's answer
2021-12-09T09:13:58-0500

The frictional torque is


"\tau_f=I\alpha=\frac12mr^2·\frac{2\pi n} t=0.0666\text{ N}·\text m."